Write a formula that expresses the relationship described below.
P varies jointly as x and y and P = 7 when x = –1 and y = –6.


I'm not sure, but I think it would be P = 6/7xy
: Write a formula that expresses the relationship described below.
P varies jointly as x and y and P = 7 when x = –1 and y = –6.


I'm not sure, but I think it would be P = 6/7xy

Answer by mangopeeler07(445)

P = 6/7xy
plug in P, x and y

7=6/7(-1)(-6)

7=6/42
-----------------------------------------------------------------------
Try one of these:

P= -(x+y)

P = 7 when x = –1 and y = –6.

7=-(-1+-6)
7=-(-7)
7=7

or the latter:

P=(7/6)(xy/1) (which perhaps is what you meant from the start)

7=7/6 times 6
7=7